OpenStudy (anonymous):

Given the equation 4x + 2y = 9, which equation below would cause an inconsistent-independent system? 6x + 12y = 13 2x + y = -6 x - 2y = 7 8x - 4y = -5 @agent0smith @kropot72 @Nnesha

Nnesha (nnesha):

x - 2y = 7 8x - 4y = -5 AND 6x + 12y = 13 2x + y = -6 like this ??

Nnesha (nnesha):

inconsistent mean two lines are parallel ( parallel lines have same slope) and there is NO SOLUTION

OpenStudy (anonymous):

SO WATS THE ANSWER

Nnesha (nnesha):

graph the lines and see which two e equations are parallel you can solve it but easy way to graph

Nnesha (nnesha):

https://www.desmos.com/calculator use this calculator and see which two lines are parallel :) and let me know what you get

Nnesha (nnesha):

wait they said independent

Nnesha (nnesha):

is it inconsistent independent or consistent independent ???

OpenStudy (anonymous):

iconsistent indepemdent

Nnesha (nnesha):

Inconsistent: mean when two lines are parallel (which mean no solution) Consistent: when lines intersect on one point and its exactly one solution Coincident: when two lines lie on top of each other (both equations are same and same line on the graph ) then you have infinite number of solutions dependent when you get exactly same line( both equations have same slope and y-intercept) independent when two lines intersect at one point (slope should different)

Nnesha (nnesha):

so graph the equations first and let me know what you get :)

OpenStudy (anonymous):

d